OpenStudy (anonymous):

Which of the following constants can be added to x^2 - 6x to form a perfect square trinomial? A.) 6 B.) 9 C.) 36

OpenStudy (the_fizicx99):

I'm going to go out on a limb here and say you're solving for the 3rd term x^2 - 6x + ___ √x^2 = x So the coefficient of the middle term 2ab = 2(x)(b) = 2xb. 2xb = 6x /2x /2x b = 3 (x-3)^2 = x^2 -2(x)(3) - 3^2 = x^2 -6x + 9

OpenStudy (anonymous):

..?

OpenStudy (the_fizicx99):

What part didn't you get?

OpenStudy (anonymous):

Any of it?

OpenStudy (the_fizicx99):

.-. K, what exactly are you trying to solve for?

OpenStudy (anonymous):

I don't know, that's why I posted it on here.

OpenStudy (the_fizicx99):

Hm, well a perfect square trinomial is in the form of (a+b)^2 or (a-b)^2 Understand so far?

OpenStudy (anonymous):

No ma'am. I'm new to all this, so I don't get any of it.

OpenStudy (the_fizicx99):

Hmm, Well a perfect square trinomial is: \(\ (a+b)^2 = a^2 + 2(a)(b) + b^2 \) understand?

OpenStudy (the_fizicx99):

And then there's the differences of squares: \(\ (a-b)^2 = a^2 -2(a)(b) - b^2 \)